Use this information to answer the question that follow.

The following data relate to direct materials costs for February:

Materials cost per yard: standard, $1.98; actual, $2.03
Standard yards per unit: standard, 4.62 yards; actual, 5.06 yards
Units of production: 9,100

Calculate the direct materials quantity variance.

a.$8,128.12 favorable

b.$7,927.92 unfavorable

c.$7,927.92 favorable

d.$8,128.12 unfavorable

Myers Corporation has the following data related to direct materials costs for November: actual costs for 4,600 pounds of material, $5.00 and standard costs for 4,470 pounds of material at $6.10 per pound.

What is the direct materials quantity variance?

a.$793 unfavorable

b.$5,060 favorable

c.$5,060 unfavorable

d.$793 favorable

Flapjack Corporation had 7,696 actual direct labor hours at an actual rate of $12.40 per hour. Original production had been budgeted for 1,100 units, but only 956 units were actually produced. Labor standards were 7.4 hours per completed unit at a standard rate of $12.96 per hour.

The direct labor rate variance is

$4,309.76 unfavorable

$4,309.76 favorable

$7,709.02 unfavorable

$7,709.02 favorable
Answer #1
direct material quantity variance = (standard quantity-actual quantity)*standard price
standarad quantity = (9100*4.62) = 42042
actual quantity = (9100*5.06) = 46046
quantity variance = (42042-46046)*1.98 = 7927.92 U
Option "c" is correct
direct material quantity variance = (standard quantity-actual quantity)*standard price
direct material quantity variance = (4470-4600)*6.10 = 793 Unfavorable
Option "a" is correct
Direct labor rate variance = (standard rate-actual rate)*actual hour
direct labor rate variance = (12.96-12.40)*7696 = 4309.76 Favorable
Option "b" is correct

Please give a thumbs up if it is helpful & let me know if any doubt